# Syntactic Language Change in English and German: Metrics, Parsers, and Convergences

### v1

@@ Abstract
Many studies have shown that human languages tend to optimize for lower complexity and increased communication efficiency. Syntactic dependency distance, which measures the linear distance between dependent words, is often considered a key indicator of language processing difficulty and working memory load. The current paper looks at diachronic trends in syntactic language change in both English and German, using corpora of parliamentary debates from the last c. 160 years. We base our observations on five dependency parsers, including the widely used Stanford CoreNLP as well as 4 newer alternatives. Our analysis of syntactic language change goes beyond linear dependency distance and explores 15 metrics relevant to dependency distance minimization (DDM) and/or based on tree graph properties, such as the tree height and degree variance. Even though we have evidence that recent parsers trained on modern treebanks are not heavily affected by data ‘noise’ such as spelling changes and OCR errors in our historic data, we find that results of syntactic language change are sensitive to the parsers involved, which is a caution against using a single parser for evaluating syntactic language change as done in previous work. We also show that syntactic language change over the time period investigated is largely similar between English and German for the different metrics explored: only 4% of cases we examine yield opposite conclusions regarding upwards and downtrends of syntactic metrics across German and English. We also show that changes in syntactic measures seem to be more frequent at the tails of sentence length distributions. To our best knowledge, ours is the most comprehensive analysis of syntactic language using modern NLP technology in recent corpora of English and German.
@@ 1 Introduction
The recent diachronic investigations involve using dependency parsers to automatically parse sentences extracted from a corpus covering a large time span, and observing how dependency distance in the parsed outputs varies over time (e.g., Lei and Wen (2020); Juzek, Krielke, and Teich (2020); Zhang and Zhou (2023)). The main advantage of such studies over approaches which leverage human annotated treebanks (Gulordava and Merlo, 2015) is that they do not require expensive manual dependency annotations and thus allow for exploring arbitrary large diachronic corpora. However, to our best knowledge, all of them rely on a single dependency parser, mostly the the Stanford CoreNLP parser (Manning et al., 2014), which was published 9 years ago. In this work, we first inspect the parsers’ legitimacy in our historical use case, as they are mostly only trained and evaluated on modern treebanks, which cover a small recent time period and may exhibit substantially less data noise (e.g., OCR errors) compared to those diachronic corpora (§4). Subsequently, we investigate whether different parsers yield the same trends regarding language change (§6.1).
@@
Further, the majority of relevant studies are for English, with the exception that Krielke (2023) investigates DDM (a.o.) in both English and German (from 1650 to 1900). She compares the trends in scientific and general language, where the texts for general language span multiple domains, such as news and fiction. In our work, we focus on a homogeneous and directly comparable genre of of political debates in both English and German, stretching from the 1800s to beyond the 2000s.
@@ 2 Related Work
Language change has been researched for a long time along multiple dimensions (and in multiple communities), including semantic, syntactic, morphological change etc. Semantic change investigates the changes in word meaning over time. For example, Giulianelli, Del Tredici, and Fernández (2020) document semantic change in the words “boy” and “girl”: Girl used to be a term for young people of either sex, while boy described male servants; the present meaning of “girl” and “boy” were first observed in the 15th and 16th centuries, respectively. Earlier works on studying semantic change relied on methods like latent semantic analysis (Sagi, Kaufmann, and Clark, 2011) or treated semantic change detection as a diachronic word sense analysis problem (Mitra et al., 2014; Frermann and Lapata, 2016). More recent studies have leveraged static or contextualized embeddings (Bamler and Mandt, 2017; Bamman, Dyer, and Smith, 2014; Chen et al., 2023; Tahmasebi, Borin, and Jatowt, 2021; Ma, Strube, and Zhao, 2024), jointly trained across varying periods (Bamler and Mandt, 2017; Bamman, Dyer, and Smith, 2014; Di Carlo, Bianchi, and Palmonari, 2019; Dubossarsky et al., 2019; Haider and Eger, 2019), or independently trained for different periods (Hamilton, Leskovec, and Jurafsky, 2016). The latter scenario then involves remapping embeddings for different periods into one shared vector space (Hamilton, Leskovec, and Jurafsky, 2016) or inducing the second-order embeddings, which represents the words’ distance to a reference vocabulary (Eger and Mehler, 2016; Rodda, Senaldi, and Lenci, 2017). Morphological change focuses on how word formation and inflection change over time (Anderson, 2015). For example, it has been shown that in most Germanic languages, including English, case morphology has gradually declined over time (Weerman and Wit, 1999). Moscoso del Prado Martín and Brendel (2016) explore case changes in Icelandic, using historical corpora annotated with part-of-speech and other morphological information. The rules of verbal inflection have been explored for many languages. For example, more frequently used English verbs have undergone faster regularization (Lieberman et al., 2007);111Note that on the contrary, highly frequent tokens in many cases resist regularization. This is known as the conservation effect Bybee (1985). In German, for instance, lower-frequency strong verbs are historically more likely to adopt the weak inflection, which has a higher type frequency. Opposed to that, highly frequent verbs tend to preserve, and in some cases even extend, irregular inflection, as discussed e.g. in Nübling (1998). verb inflection tends to become weaker in German (Carroll, Svare, and Salmons, 2012) and Dutch (Knooihuizen and Strik, 2014) over time. Zhu and Lei (2018) further confirm inflection decay, i.e., languages become less inflectional, in Modern English, using entropy-based algorithms. Syntactic change has been discussed for changes in specific syntactic patterns and/or metrics reflecting language complexity. For instance, Krielke (2021); Krielke et al. (2022); Krielke (2023); Degaetano-Ortlieb and Teich (2022) investigate syntactic shifts in scientific English and/or German, observing several phenomena, such as the tendency towards heavy noun phrases and a simpler sentence structure with decreasing usage of clauses over time. Similar phenomena were also observed in Halliday (2003); Banks (2008); Biber and Gray (2011, 2016). One of the widely adopted approaches to model language complexity revolves around syntactic dependency distance. Studies have indicated that humans tend to position syntactically related words closer together (Tily, 2010; Gulordava and Merlo, 2015; Liu, Xu, and Liang, 2017; Zhang and Zhou, 2023); this hypothesis/phenomenon is called DDM.
@@
More recent studies leverage dependency parsers to automatically parse sentences extracted from diachronic corpora, and then base their observations on the parsing results (Lei and Wen, 2020; Juzek, Krielke, and Teich, 2020; Liu, Zhu, and Lei, 2022; Zhu, Liu, and Pang, 2022; Zhang and Zhou, 2023; Krielke, 2023). An important benefit of such approaches is that they do not require expensive dependency annotations; therefore, arbitrary corpora beyond treebanks can serve as the research resource. Lei and Wen (2020); Liu, Zhu, and Lei (2022) investigate the diachronic trend of dependency distance based on the State of the Union Addresses corpus,222https://www.presidency.ucsb.edu/ which contains the political speeches of 43 U.S. presidents from 1790 to 2017. They show that the mDD and its normalized variant (nDD) (Lei and Jockers, 2020) of American English sentences are decreasing over time. Lei and Wen (2020) also observe a consistent downward trend across different sentence length groups. Liu, Zhu, and Lei (2022) additionally divide the sentences of 10 words (the shortest sentence group in Lei and Wen (2020)) into two subgroups: sentences of 0-4 and 5-10 words, finding the existence of an anti-DDM phenomenon for the 0-4 group, i.e., DDM does not apply to short sentences because of the constraints by other language optimization principles (Ferrer-i Cancho and Gómez-Rodríguez, 2021). Zhu, Liu, and Pang (2022); Zhang and Zhou (2023) compare the diachronic trend of mDD and/or nDD for the past 100-200 years across varying text genres, based on the COHA corpus,333https://www.english-corpora.org/coha/ which contains texts in American English across 4 domains: News, Magazine, Fiction, Non-fiction. Their findings are often inconsistent. E.g., Zhu, Liu, and Pang (2022) find a downtrend of mDD only for the Fiction genre but an upward trend for the News and Magazine genres during the period 1900-2009. In contrast, Zhang and Zhou (2023) observe a continuous decrease in mDD and nDD for the Magazine genre spanning from 1815 to 2009. Nevertheless, both of their results imply that the diachronic trends of mDD vary for different genres. Juzek, Krielke, and Teich (2020) explore diachronic shifts in syntactic patterns for scientific English, leveraging the Royal Society Corpus (Fischer et al., 2020), which contains scientific publications from 1665 to 1996. They also observe a decrease in mDD over time. As follow-up work, Krielke (2023) comprehensively investigates language optimization for scientific English and German from 1650 to 1900, in comparison to that for general language. She finds that while the mDD for scientific language decreased over time, mDD for general-domain language does not show a decrease over time during the same period.
@@
However, all those approaches have three main limitations: (1) all of them rely on one single parser, mostly the Stanford CoreNLP parser (Manning et al., 2014) with an exception that Krielke (2023) leverages the UDPipe parser (Straka, 2018). In our work, we explore whether different parsers yield the same outcomes regarding our diachronic investigations. (2) Some of the works may report biased results due to improper sentence grouping. We show a clear trend of shorter sentences appearing more frequently in later periods, consequently, the average sentence length per decade is decreasing over time. Zhang and Zhou (2023) do not differentiate sentences by lengths and report the diachronic trends on all sentences. Liu, Zhu, and Lei (2022); Lei and Wen (2020) divide the sentences into groups having 1-10 (0-4, 5-10), 11-20, 21-30, and 31+ words, and analyze trends for each group separately. However, we demonstrate that even when restricting the maximum difference in sentence lengths to 3 tokens within a group, a slight downtrend in sentence length over time persists (refer to §7.0.1). Given the substantial length disparity within a group in their works (e.g., 10 for group 11-20), and lacking information on the length distribution in their data, it is hard to tell whether the observed downtrend of mDD is the effect of decreasing sentence length or really due to DDM. Those facts cast our doubts on the legitimacy of their findings. (3) Almost all of the works focus on English change, except for Krielke (2023), who compares the changes in scientific and general English and German. The dataset for general languages used in Krielke (2023) contain texts in multiple genres like fiction and news. Nevertheless, as Zhu, Liu, and Pang (2022); Zhang and Zhou (2023) suggest, trends of mDD may differ for various text genres. Instead, in this work, we focus on one specific genre, namely political debates, across both languages.
@@ 3 Data Preparation
In this work, we use corpora consisting of political debates and speeches in German and English to ensure a fair comparison. For English, we select the Hansard corpus, which comprises the official reports of UK parliament debates since 1803. To compile the data for the time period 1803—2004, we extract XML files from the Hansard archive,444https://www.hansard-archive.parliament.uk/ retaining every section labeled as “debate”. Data from 2005 to April 2021 is obtained from a Zenodo repository.555https://evanodell.com/projects/datasets/hansard-data/ (v3.1.0) For German, we utilize the DeuParl corpus published by Beese, Pütz, and Eger (2022); it contains plenary protocols from both the Reichstag (the former German parliament, until 1945) and the Bundestag (the current German parliament), covering the period from 1867 to 2022.
@@
“The Earl of Liverpool […] [238 stance, when he […] |
Blankets 18,800 13,500 6,360 |
Ich will Ihnen z. B. sagen, ist […] |
(Bravo! Rechts.) Reichstag des […] 20. Sitzung am 27. März 1867. |
@@ 3.1 Validation & Correction
To validate the feasibility of our preprocessing pipeline and study the influence of data noise on metrics such as dependency distance (discussed in Section 6), for each corpus, we apply the pipeline to five randomly sampled paragraphs from each decade, resulting in approximately 150-800 sentences per decade. We then manually review ten randomly selected outputs for each decade and make corrections if any issues are identified. Additionally, the data from the 2020s is merged into the 2010s, given that only partial data is available for the former.
@@
We first check the German sentences from DeuParl. Two annotators with an NLP background, one male faculty member who is a German native, and one female PhD student who speaks German fluently, check the outputs. In this phase, we do not have annotation guidelines; the annotators are asked to freely make comments on each text following their intuition and make corrections if they identify any issues in the text.666The annotation is done with Google spreadsheet. Subsequently, by inspecting the comments and comparing them to the original PDF files,777https://www.reichstagsprotokolle.de/ we standardize and format the annotation scheme. As Table 2 shows, the formatted annotation has 5 columns:
| text | date | is_sent | has_issues | issues | origin_of_issues | correction |
Ich bitte, daß diejenigen Herren, welche für den Fall der Annahme des Z 33b in demselben die Worte „oder an Druck m anderen öffentlichen Orten" aufrecht erhalten wollen, sich von ihren Plätzen erheben. |
1883-4-6 |
TRUE |
TRUE |
extral material; symbol; spelling |
ocr;ocr;historic |
Ich bitte, dass diejenigen Herren, welche für den Fall der Annahme des § 33b in demselben die Worte „oder an anderen öffentlichen Orten" aufrecht erhalten wollen, sich von ihren Plätzen erheben. |
@@
(2) has_issues: A binary annotation about the existence of any issues (only if (1) is ‘TRUE’).
@@
Category |
Example |
Correction |
Spelling |
Das ist die Mehrheit; die Diskussion ist geschloffen. |
[…] geschlossen. |
Space |
Der EVG-Vertrag spreche von der „westlichen Verteidigung", und im Protokoll der NATO-Staaten sei vom Zusammenschluß der w e s t europäischen Länder die Rede. |
[…] westeuropäischen […] |
Missing Material |
Ich bin der Meinung — und viele an uns herangekommene Klagen lassen auch darauf schließen — , dass die Auffassung des Reichssparkommissars, dass das Personal der Deutschen Reichspost voll ausgelastet, aber nicht überlastet sei, nicht richtig ist. |
add “—” |
Extra Material |
[…] daß durch den Bund zweierlei Recht für die Norddeutschen geschaffen werden soll, (Sehr richtig!) daß gewissermaßen zweierlei Klassen von Norddeutschen geschaffen werden sollen, (Sehr gut!) eine Selekta, die vermöge ihrer Gesittung […] |
delete “(Sehr gut!)” and “(Sehr richtig!)” |
Punctuation & Symbol |
Ich bitte, daß diejenigen Herren, welche für den Fall der Annahme des Z 33b in demselben die Worte „oder an Druck m anderen öffentlichen Orten" aufrecht erhalten wollen, sich von ihren Plätzen erheben. |
[…] § 33b […] |
We classify the issues into 5 categories: Spelling, Space, Missing Material, Extra Material and Punctuation & Symbol, as shown in Table 3. Additionally, we split the origins of issues into 4 categories: (1) Historic, which indicates the issues are due to the differences in historical and modern language. For example, “daß” is a spelling issue with the historic origin. (2) OCR, which indicates the issue is an OCR error. For example, the spelling issue in Table 3 results from an OCR error (the archaic long S in word medial or initial position is misanalysed as ⟨f⟩). (3) Genre, which represents issues that exist due to the characteristics of the text genre. In the text of political debates, the language changes between written and spoken language frequently, and also, there exist interjections within a sentence, as shown in the fourth row of Table 3 (extra material). (4) Preprocessing, which denotes that the issues stem from our preprocessing pipeline, e.g., extra space. For details regarding those categories, we refer to Figure 21 in the appendix, which demonstrates our annotation guidelines following the refined annotation scheme.
@@
To check the agreement among annotators, 40 German sentences are jointly annotated, 20 for the sentence identification task and 20 for the issue identification/correction task. For English, 30 sentences are commonly annotated for all annotation subtasks. For each binary annotation task, we calculate Cohen’s Kappa for inter-agreement, whereas for the correction task, we compute the edit distance between sentences. Let be the original sentence, and and be the sentences corrected by annotators A and B respectively. We calculate the edit distance (i) between and (), (ii) between and (), and (iii) between and (). If the corrections are similar to each other, (i) and (ii) should be close while (iii) should be small. Overall, the annotators obtained a decent level of agreement, with 0.7 on binary tasks and the edit distances of (i) = 3.88, (ii) = 3.96, and (iii) = 0.54 for the correction.

### v2

@@ Abstract
Many studies have shown that human languages tend to optimize for lower complexity and increased communication efficiency. Syntactic dependency distance, which measures the linear distance between dependent words, is often considered a key indicator of language processing difficulty and working memory load. The current paper looks at diachronic trends in syntactic language change in both English and German, using corpora of parliamentary debates from the last c. 160 years. We base our observations on five dependency parsers, including the widely used Stanford CoreNLP as well as 4 newer alternatives. Our analysis of syntactic language change goes beyond linear dependency distance and explores 15 metrics relevant to dependency distance minimization (DDM) and/or based on tree graph properties, such as the tree height and degree variance. Even though we have evidence that recent parsers trained on modern treebanks are not heavily affected by data ‘noise’ such as spelling changes and OCR errors in our historic data, we find that results of syntactic language change are sensitive to the parsers involved, which is a caution against using a single parser for evaluating syntactic language change as done in previous work. We also show that syntactic language change over the time period investigated is largely similar between English and German for the different metrics explored: only 4% of cases we examine yield opposite conclusions regarding upwards and downtrends of syntactic metrics across German and English. We also show that changes in syntactic measures seem to be more frequent at the tails of sentence length distributions. To our best knowledge, ours is the most comprehensive analysis of syntactic language change using modern NLP technology in recent corpora of English and German.
@@ 1 Introduction
The recent diachronic investigations involve using dependency parsers to automatically parse sentences extracted from a corpus covering a large time span, and observing how dependency distance in the parsed outputs varies over time (e.g., Lei and Wen (2020); Juzek, Krielke, and Teich (2020); Zhang and Zhou (2023)). The main advantage of such studies over approaches which leverage human annotated treebanks (Gulordava and Merlo, 2015) is that they do not require expensive manual dependency annotations and thus allow for exploring arbitrary large diachronic corpora. However, to our best knowledge, all of them rely on a single dependency parser, mostly the the Stanford CoreNLP parser (Manning et al., 2014), which was published 9 years ago. In this work, we first inspect the parsers’ legitimacy in our historical use case, as they are mostly only trained and evaluated on modern treebanks, which cover a small recent time period and may exhibit substantially less data noise (e.g., OCR errors) compared to historical corpora (§4). Subsequently, we investigate whether different parsers yield the same trends regarding language change (§6.1).
@@
Further, the majority of relevant studies are for English, with the exception that Krielke (2023) investigates DDM (a.o.) in both English and German (from 1650 to 1900). She compares the trends in scientific and general language, where the texts for general language span multiple domains, such as news and fiction. In our work, we focus on a homogeneous and directly comparable genre of political debates in both English and German, stretching from the 1800s to beyond the 2000s.
@@
We will make our code+data public upon acceptance.111https://github.com/cyr19/syntaxchange
@@ 2 Related Work
Language change has been researched for a long time along multiple dimensions (and in multiple communities), including semantic, syntactic, morphological change etc. Semantic change investigates the changes in word meaning over time. For example, Giulianelli, Del Tredici, and Fernández (2020) document semantic change in the words “boy” and “girl”: Girl used to be a term for young people of either sex, while boy described male servants; the present meaning of “girl” and “boy” were first observed in the 15th and 16th centuries, respectively. Earlier works on studying semantic change relied on methods like latent semantic analysis (Sagi, Kaufmann, and Clark, 2011) or treated semantic change detection as a diachronic word sense analysis problem (Mitra et al., 2014; Frermann and Lapata, 2016). More recent studies have leveraged static or contextualized embeddings (Bamler and Mandt, 2017; Bamman, Dyer, and Smith, 2014; Chen et al., 2023; Tahmasebi, Borin, and Jatowt, 2021; Rother, Haider, and Eger, 2020; Ma, Strube, and Zhao, 2024), jointly trained across varying periods (Bamler and Mandt, 2017; Bamman, Dyer, and Smith, 2014; Di Carlo, Bianchi, and Palmonari, 2019; Dubossarsky et al., 2019; Haider and Eger, 2019), or independently trained for different periods (Hamilton, Leskovec, and Jurafsky, 2016). The latter scenario then involves remapping embeddings for different periods into one shared vector space (Hamilton, Leskovec, and Jurafsky, 2016) or inducing the second-order embeddings, which represents the words’ distance to a reference vocabulary (Eger and Mehler, 2016; Rodda, Senaldi, and Lenci, 2017). Morphological change focuses on how word formation and inflection change over time (Anderson, 2015). For example, it has been shown that in most Germanic languages, including English, case morphology has gradually declined over time (Weerman and Wit, 1999). Moscoso del Prado Martín and Brendel (2016) explore case changes in Icelandic, using historical corpora annotated with part-of-speech and other morphological information. The rules of verbal inflection have been explored for many languages. For example, more frequently used English verbs have undergone faster regularization (Lieberman et al., 2007);222Note that on the contrary, highly frequent tokens in many cases resist regularization. This is known as the conservation effect Bybee (1985). In German, for instance, lower-frequency strong verbs are historically more likely to adopt the weak inflection, which has a higher type frequency. Opposed to that, highly frequent verbs tend to preserve, and in some cases even extend, irregular inflection, as discussed e.g. in Nübling (1998). verb inflection tends to become weaker in German (Carroll, Svare, and Salmons, 2012) and Dutch (Knooihuizen and Strik, 2014) over time. Zhu and Lei (2018) further confirm inflection decay, i.e., languages become less inflectional, in Modern English, using entropy-based algorithms. Syntactic change has been discussed for changes in specific syntactic patterns and/or metrics reflecting language complexity. For instance, Krielke (2021); Krielke et al. (2022); Krielke (2023); Degaetano-Ortlieb and Teich (2022) investigate syntactic shifts in scientific English and/or German, observing several phenomena, such as the tendency towards heavy noun phrases and a simpler sentence structure with decreasing usage of clauses over time. Similar phenomena were also observed in Halliday (2003); Banks (2008); Biber and Gray (2011, 2016). One of the widely adopted approaches to model language complexity revolves around syntactic dependency distance. Studies have indicated that humans tend to position syntactically related words closer together (Tily, 2010; Gulordava and Merlo, 2015; Liu, Xu, and Liang, 2017; Zhang and Zhou, 2023); this hypothesis/phenomenon is called DDM.
@@
More recent studies leverage dependency parsers to automatically parse sentences extracted from diachronic corpora, and then base their observations on the parsing results (Lei and Wen, 2020; Juzek, Krielke, and Teich, 2020; Liu, Zhu, and Lei, 2022; Zhu, Liu, and Pang, 2022; Zhang and Zhou, 2023; Krielke, 2023). An important benefit of such approaches is that they do not require expensive dependency annotations; therefore, arbitrary corpora beyond treebanks can serve as the research resource. Lei and Wen (2020); Liu, Zhu, and Lei (2022) investigate the diachronic trend of dependency distance based on the State of the Union Addresses corpus,333https://www.presidency.ucsb.edu/ which contains the political speeches of 43 U.S. presidents from 1790 to 2017. They show that the mDD and its normalized variant (nDD) (Lei and Jockers, 2020) of American English sentences are decreasing over time. While Lei and Wen (2020) observe a consistent downward trend across different sentence length groups, Liu, Zhu, and Lei (2022) additionally divide the sentences of 10 words (the shortest sentence group in Lei and Wen (2020)) into two subgroups: sentences of 0-4 and 5-10 words, finding the existence of an anti-DDM phenomenon for the 0-4 group, i.e., DDM does not apply to short sentences because of the constraints by other language optimization principles (Ferrer-i Cancho and Gómez-Rodríguez, 2021). Zhu, Liu, and Pang (2022); Zhang and Zhou (2023) compare the diachronic trend of mDD and/or nDD for the past 100-200 years across varying text genres, based on the COHA corpus,444https://www.english-corpora.org/coha/ which contains texts in American English across 4 domains: News, Magazine, Fiction, Non-fiction. Their findings are often inconsistent. E.g., Zhu, Liu, and Pang (2022) find a downtrend of mDD only for the Fiction genre but an upward trend for the News and Magazine genres during the period 1900-2009. In contrast, Zhang and Zhou (2023) observe a continuous decrease in mDD and nDD for the Magazine genre spanning from 1815 to 2009. Nevertheless, both of their results imply that the diachronic trends of mDD vary for different genres. Juzek, Krielke, and Teich (2020) explore diachronic shifts in syntactic patterns for scientific English, leveraging the Royal Society Corpus (Fischer et al., 2020), which contains scientific publications from 1665 to 1996. They also observe a decrease in mDD over time. As follow-up work, Krielke (2023) comprehensively investigates language optimization for scientific English and German from 1650 to 1900, in comparison to that for general language. She finds that while the mDD for scientific language decreased over time, mDD for general-domain language does not show a decrease over time during the same period.
@@
However, all those approaches have three main limitations: (1) all of them rely on one single parser, mostly the Stanford CoreNLP parser (Manning et al., 2014) with an exception that Krielke (2023) leverages the UDPipe parser (Straka, 2018). In our work, we explore whether different parsers yield the same outcomes regarding our diachronic investigations. (2) Some of the works may report biased results due to improper sentence grouping. We show a clear trend of shorter sentences appearing more frequently in later periods, consequently, the average sentence length per decade is decreasing over time. Zhang and Zhou (2023) do not differentiate sentences by lengths and report the diachronic trends on all sentences. Liu, Zhu, and Lei (2022); Lei and Wen (2020) divide the sentences into groups having 1-10 (0-4, 5-10), 11-20, 21-30, and 31+ words, and analyze trends for each group separately. However, we demonstrate that even when restricting the maximum difference in sentence lengths to 3 tokens within a group, a slight downtrend in sentence length over time persists (refer to §7.0.1). Given the substantial length disparity within a group in their works (e.g., 10 for group 11-20), and lacking information on the length distribution in their data, it is hard to tell whether the observed downtrend of mDD is the effect of decreasing sentence length or really due to DDM. Those facts may cast doubts on the legitimacy of their findings. (3) Almost all of the works focus on English change, except for Krielke (2023), who compares the changes in scientific and general English and German. The dataset for general languages used in Krielke (2023) contain texts in multiple genres like fiction and news. Nevertheless, as Zhu, Liu, and Pang (2022); Zhang and Zhou (2023) suggest, trends of mDD may differ for various text genres. Instead, in this work, we focus on one specific genre, namely political debates, across both languages.
@@ 3 Data Preparation
In this work, we use corpora consisting of political debates and speeches in German and English to ensure a fair comparison. For English, we select the Hansard corpus, which comprises the official reports of UK parliament debates since 1803. To compile the data for the time period 1803—2004, we extract XML files from the Hansard archive,555https://www.hansard-archive.parliament.uk/ retaining every section labeled as “debate”. Data from 2005 to April 2021 is obtained from a Zenodo repository.666https://evanodell.com/projects/datasets/hansard-data/ (v3.1.0) For German, we utilize the DeuParl corpus published by Beese, Pütz, and Eger (2022);777The original corpus was published in Walter et al. (2021). The version we use contains further clean-up. it contains plenary protocols from both the Reichstag (the former German parliament, until 1945) and the Bundestag (the current German parliament), covering the period from 1867 to 2022.
@@
| “The Earl of Liverpool […] [238 stance, when he […] | Blankets 18,800 13,500 6,360 | Ich will Ihnen z. B. sagen, ist […] | (Bravo! Rechts.) Reichstag des […] 20. Sitzung am 27. März 1867. |
@@ 3.1 Validation & Correction
To validate the feasibility of our preprocessing pipeline and study the influence of data noise on metrics such as dependency distance (discussed in §5.1), for each corpus, we apply the pipeline to five randomly sampled paragraphs from each decade, resulting in approximately 150-800 sentences per decade. We then manually review ten randomly selected outputs for each decade and make corrections if any issues are identified. Additionally, the data from the 2020s is merged into the 2010s, given that only partial data is available for the former.
| text | date | is sent | has issues | issues | origin of issues | correction |
| Ich bitte, daß diejenigen Herren, welche für den Fall der Annahme des Z 33b in demselben die Worte „oder an Druck m anderen öffentlichen Orten" aufrecht erhalten wollen, sich von ihren Plätzen erheben. | 1883-4-6 | TRUE | TRUE | extra material; symbol; spelling | ocr; ocr; historic | Ich bitte, dass diejenigen Herren, welche für den Fall der Annahme des § 33b in demselben die Worte „oder an anderen öffentlichen Orten" aufrecht erhalten wollen, sich von ihren Plätzen erheben. |
@@
We first check the German sentences from DeuParl. Two annotators with an NLP background, one male faculty member who is a German native, and one female PhD student who speaks German fluently, check the outputs. In this phase, we do not have annotation guidelines; the annotators are asked to freely make comments on each text following their intuition and make corrections if they identify any issues in the text.888The annotation is done with Google spreadsheet. Subsequently, by inspecting the comments and comparing them to the original PDF files,999https://www.reichstagsprotokolle.de/ we standardize and format the annotation scheme. As Table 2 shows, the formatted annotation has 5 columns:
@@
(2) has_issues: A binary annotation (‘TRUE’/‘FALSE’) about the existence of any issues (only if (1) is ‘TRUE’).
@@
| Category | Example | Correction |
| Spelling | Das ist die Mehrheit; die Diskussion ist geschloffen. | […] geschlossen. |
| Space | Der EVG-Vertrag spreche von der „westlichen Verteidigung", und im Protokoll der NATO-Staaten sei vom Zusammenschluß der w e s t europäischen Länder die Rede. | […] westeuropäischen […] |
| Missing Material | Ich bin der Meinung — und viele an uns herangekommene Klagen lassen auch darauf schließen — , dass die Auffassung des Reichssparkommissars, dass das Personal der Deutschen Reichspost voll ausgelastet, aber nicht überlastet sei, nicht richtig ist. | add “—” |
| Extra Material | […] daß durch den Bund zweierlei Recht für die Norddeutschen geschaffen werden soll, (Sehr richtig!) daß gewissermaßen zweierlei Klassen von Norddeutschen geschaffen werden sollen, (Sehr gut!) eine Selekta, die vermöge ihrer Gesittung […] | delete “(Sehr gut!)” and “(Sehr richtig!)” |
| Punctuation & Symbol | Ich bitte, daß diejenigen Herren, welche für den Fall der Annahme des Z 33b in demselben die Worte „oder an Druck m anderen öffentlichen Orten" aufrecht erhalten wollen, sich von ihren Plätzen erheben. | […] § 33b […] |
We classify the issues into 5 categories: Spelling, Space, Missing Material, Extra Material and Punctuation & Symbol, as shown in Table 3. Additionally, we split the origins of issues into 4 categories: (1) Historic, which indicates the issues are due to the differences in historical and modern language. For example, ‘daß’ (modern spelling: ‘dass’) is a spelling issue with the historic origin. (2) OCR, which indicates the issue is an OCR error. For example, the spelling issue in Table 3 results from an OCR error (‘geschloffen’ vs. ‘geschlossen’). (3) Genre, which represents issues that exist due to the characteristics of the text genre. In the text of political debates, the language changes between written and spoken language frequently, and also, there exist interjections within a sentence, as shown in the fourth row of Table 3 (Extra Material). (4) Preprocessing, which denotes that the issues stem from our preprocessing pipeline, e.g., extra space. For details regarding those categories, we refer to Figure 18 in the appendix, which demonstrates our annotation guidelines following the refined annotation scheme.
@@
To check the agreement among annotators, 40 German sentences are jointly annotated, 20 for the sentence identification task and 20 for the issue identification/correction task. For English, 30 sentences are commonly annotated for all annotation subtasks. For each binary annotation task, we calculate Cohen’s Kappa (Cohen, 1960) for inter-agreement, whereas for the correction task, we compute the edit distance (Levenshtein, 1965) between sentences. Let be the original sentence, and and be the sentences corrected by annotators A and B respectively. We calculate the edit distance (i) between and (), (ii) between and (), and (iii) between and (). If the corrections are similar to each other, (i) and (ii) should be close while (iii) should be small. Overall, the annotators obtained a decent level of agreement, with 0.7 on binary tasks and the edit distances of (i) = 3.88, (ii) = 3.96, and (iii) = 0.54 for the correction.